A motorboat goes 5 miles upstream in the same time it requires to 7 miles downstream. If the river flows at 2mph, find the speed of the boat in still water.
Answer by mananth(12270)

let the speed of boat be x mph
speed of current = 2mph
..
upstream speed = x-2 mph
downstream speed = x+2 mph
..
Time = distance/speed
time upstream time downstrea
..
5/(x-2)=7/(x+2)
cross multiply
5(x+2)=7(x-2)
5x+10=7x-14
2x=24
/2
x=12 mph====> speed of boat